# Scaling Transformers for Discriminative Recommendation via Generative Pretraining

### v1

@@ Abstract.
Discriminative recommendation tasks, such as CTR (click-through rate) and CVR (conversion rate) prediction, play critical roles in the ranking stage of large-scale industrial recommender systems. However, training a discriminative model encounters a significant overfitting issue induced by data sparsity. Moreover, this overfitting issue worsens with larger models, causing them to underperform smaller ones. To address the overfitting issue and enhance model scalability, we propose a framework named GPSD (Generative Pretraining for Scalable Discriminative Recommendation), drawing inspiration from generative training, which exhibits no evident signs of overfitting. GPSD leverages the parameters learned from a pretrained generative model to initialize a discriminative model, and subsequently applies a sparse parameter freezing strategy. Extensive experiments conducted on both industrial-scale and publicly available datasets demonstrate the superior performance of GPSD. Moreover, it delivers remarkable improvements in online A/B tests. GPSD offers two primary advantages: 1) it substantially narrows the generalization gap in model training, resulting in better test performance; and 2) it leverages the scalability of Transformers, delivering consistent performance gains as models are scaled up. Specifically, we observe consistent performance improvements as the model dense parameters scale from 13K to 0.3B, closely adhering to power laws. These findings pave the way for unifying the architectures of recommendation models and language models, enabling the direct application of techniques well-established in large language models to recommendation models.
@@
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15567745.
@@ 3.3. Revisiting the Overfitting Phenomenon
Figure 3(b) shows the training and validation AUC curves of discriminative models trained for the CTR task. Regardless of model scale, a significant generalization gap exists between training and validation performance, indicating severe overfitting. Consequently, while the larger model (L4H256A4) achieves significantly better training performance, it achieve worse validation performance than the smaller model (L4H128A4). Specifically, we identify two distinct types of overfitting. As shown in Figure 3(b), models across four scales all exhibit abrupt performance drop at epoch transitions. This phenomenon, referred to as one-epoch overfitting, has been investigated in (Zhang et al., 2022). We confirm the occurrence of this phenomenon on a industrial scale dataset, which is 16×16$\times$16 × the size of the largest dataset used in (Zhang et al., 2022). Besides one-epoch overfitting, there is a second type of overfitting, which is subtler. After certain steps within the first epoch, the validation AUC almost stagnates while the training AUC continues to grow fast. We name this type of overfitting within-one-epoch overfitting in order to correspond with one-epoch overfitting. Both overfitting phenomena hinder the scalability of Transformers for discriminative tasks and obstruct the path of replicating the successes of large language models achieved by scaling.
In contrast to discriminative models that face severe overfitting, we find generative models demonstrate robustness against this issue. As shown in Figure 3(a), the training and validation loss curves of generative models remain a small constant gap throughout the training process. The constant generalization gap is expected and acceptable, often caused by distributional shifts occurring over time. This inherent resistance to overfitting leads to improved scalability, with larger models consistently achieve superior performance compared to their smaller counterparts. We hypothesize that generative training avoids the sparsity issue through extensive random negative sampling, thus leading to more stable and sufficient training of sparse parameters.
(a) Loss curves of generative models.
(b) AUC curves of discriminative models for the CTR task.
@@ 3.4. Enhancing Discriminative Training via Generative Pretraining
Figure 3 demonstrates that generative models, unlike discriminative models, do not face severe overfitting problem and can achieve better performance when scaled up. This discrepancy inspires us to enhance discriminative training via generative pretraining. As introduced in Section 2.3, there are various strategies to bridge a pretrained generative model and a discriminative model. In this section, we conduct experiments on these strategies and try to figure out which one is better. The results are shown in Table 3. According to results, we can draw the following conclusions:
@@
To further demonstrate why the SF strategy is effective, we present the AUC curves in Figure 4. Figures 4(a) and 4(b) reveal that without the SF strategy, the benefits of pretraining are limited, and the model suffers from the same overfitting issue as observed in the baseline model (Figure 3(b)). Figures 4(c) and 4(d) reveal that the SF strategy successfully addresses both one-epoch and within-one-epoch overfitting phenomena while substantially narrowing the generalization gap, leading to significantly better test performance.
@@ 3.6. Scaling Up Further
We leverage the ST&SF strategy in this experiment due to its flexibility of decoupling the architecture of generative pretraining and discriminative training. The model settings used are listed in Table 5. For the pretraining phase, we use networks fixed at 4 layers with varying widths. This reduces resource consumption while maintaining comparable downstream performance. For discriminative training, we apply a smaller peak learning rate for larger models for better training stability. We scale the sparse parameters from 125M to 4B (32×32$\times$32 ×) and scale the dense parameters from 13K to 327M ($25K\times$).
@@ 3.7. Cross-Architecture Transfer
xx
@@ 3.9. Online A/B Results
We apply the GPSD framework to the ranking model in the product recommender system of e-commerce platform AliExpress [^3] https://www.aliexpress.com. The base ranking model is a feature-rich multitask model, with hundreds of categorical and numerical features and three tasks. It employs single-layer target attention modules to encode user behavior items and is incrementally trained every day on new comming data. To apply the GPSD framework onto the base model, we develop an incremental-GPSD framework that integrates incremental training with the ST&SF strategy, as illustrated in Figure 7. We also substitute the target attention modules with Transformers to encode user behavior items. For online efficiency, we only employ small-scale Transformers, i.e., L3H160A4. Despite the small scale, the model still achieves significant online gains, with GMV increasing by 7.97% and the number of buyers increasing by 1.79%. These results are based on a 7-day experimental window.

### v2

@@ Abstract.
Discriminative recommendation tasks, such as CTR (click-through rate) and CVR (conversion rate) prediction, play critical roles in the ranking stage of large-scale industrial recommender systems. However, training a discriminative model encounters a significant overfitting issue induced by data sparsity. Moreover, this overfitting issue worsens with larger models, causing them to underperform smaller ones. To address the overfitting issue and enhance model scalability, we propose a framework named GPSD (Generative Pretraining for Scalable Discriminative Recommendation), drawing inspiration from generative training, which exhibits no evident signs of overfitting. GPSD leverages the parameters learned from a pretrained generative model to initialize a discriminative model, and subsequently applies a sparse parameter freezing strategy. Extensive experiments conducted on both industrial-scale and publicly available datasets demonstrate the superior performance of GPSD. Moreover, it delivers remarkable improvements in online A/B tests. GPSD offers two primary advantages: 1) it substantially narrows the generalization gap in model training, resulting in better test performance; and 2) it leverages the scalability of Transformers, delivering consistent performance gains as models are scaled up. Specifically, we observe consistent performance improvements as the model dense parameters scale from 13K to 0.3B, closely adhering to power laws. These findings pave the way for unifying the architectures of recommendation models and language models, enabling the direct application of techniques well-established in large language models to recommendation models. The code is available at https://github.com/chqiwang/gpsd-rec.
@@ 3.3. Revisiting the Overfitting Phenomenon
Figure 3(a) shows the training and validation AUC curves of discriminative models trained for the CTR task. Regardless of model scale, a significant generalization gap exists between training and validation performance, indicating severe overfitting. Consequently, while the larger model (L4H256A4) achieves significantly better training performance, it achieve worse validation performance than the smaller model (L4H128A4). Specifically, we identify two distinct types of overfitting. As shown in Figure 3(a), models across four scales all exhibit abrupt performance drop at epoch transitions. This phenomenon, referred to as one-epoch overfitting, has been investigated in (Zhang et al., 2022). We confirm the occurrence of this phenomenon on a industrial scale dataset, which is $16\times$ the size of the largest dataset used in (Zhang et al., 2022). Besides one-epoch overfitting, there is a second type of overfitting, which is subtler. After certain steps within the first epoch, the validation AUC almost stagnates while the training AUC continues to grow fast. We name this type of overfitting within-one-epoch overfitting in order to correspond with one-epoch overfitting. Both overfitting phenomena hinder the scalability of Transformers for discriminative tasks and obstruct the path of replicating the successes of large language models achieved by scaling.
In contrast to discriminative models that face severe overfitting, we find generative models demonstrate robustness against this issue. As shown in Figure 3(b), the training and validation loss curves of generative models remain a small constant gap throughout the training process. The constant generalization gap is expected and acceptable, often caused by distributional shifts occurring over time. This inherent resistance to overfitting leads to improved scalability, with larger models consistently achieve superior performance compared to their smaller counterparts. We hypothesize that generative training avoids the sparsity issue through extensive random negative sampling, thus leading to more stable and sufficient training of sparse parameters.
(a) AUC curves of discriminative models for the CTR task.
(b) Loss curves of generative models.
@@
A comparison of training behaviors between generative and discriminative models.
@@ 3.4. Enhancing Discriminative Training via Generative Pretraining
Figure 3(b) demonstrates that generative models, unlike discriminative models, do not face severe overfitting problem and can achieve better performance when scaled up. This discrepancy inspires us to enhance discriminative training via generative pretraining. As introduced in Section 2.3, there are various strategies to bridge a pretrained generative model and a discriminative model. In this section, we conduct experiments on these strategies and try to figure out which one is better. The results are shown in Table 3. According to results, we can draw the following conclusions:
@@
To further demonstrate why the SF strategy is effective, we present the AUC curves in Figure 4. Figures 4(a) and 4(b) reveal that without the SF strategy, the benefits of pretraining are limited, and the model suffers from the same overfitting issue as observed in the baseline model (Figure 3(a)). Figures 4(c) and 4(d) reveal that the SF strategy successfully addresses both one-epoch and within-one-epoch overfitting phenomena while substantially narrowing the generalization gap, leading to significantly better test performance.
@@
AUC curves of discriminative models for the CTR task under various strategies.
@@ 3.6. Scaling Up Further
We leverage the ST&SF strategy in this experiment due to its flexibility of decoupling the architecture of generative pretraining and discriminative training. The model settings used are listed in Table 5. For the pretraining phase, we use networks fixed at 4 layers with varying widths. This reduces resource consumption while maintaining comparable downstream performance. For discriminative training, we apply a smaller peak learning rate for larger models for better training stability. We scale the sparse parameters from 125M to 4B ($32\times$) and scale the dense parameters from 13K to 327M ($25K\times$).
@@ 3.7. Cross-Architecture Transfer
Results of cross-architecture transfer from Transformer to HSTU and Wukong.
@@ 3.9. Online A/B Results
We apply the GPSD framework to the ranking model in the product recommender system of e-commerce platform AliExpress [^3] https://www.aliexpress.com. The base ranking model is a feature-rich multitask model, with hundreds of categorical and numerical features and three tasks. It employs single-layer target attention modules to encode user behavior items and is incrementally trained every day on new comming data. To apply the GPSD framework onto the base model, we develop an incremental-GPSD framework that integrates incremental training with the ST&SF strategy, as illustrated in Figure 7. We also substitute the target attention modules with Transformers to encode user behavior items. For online efficiency, we only employ small-scale Transformers, i.e., L3H160A4. Despite the small scale, the model still achieves significant online gains as shown in Table 7.
@@
Illustration of the incremental-GPSD framework. Both the generative model and the ranking model are trained in an incremental manner. In each iteration of the ranking model training, the sparse parameters from the generative model are first transferred to the ranking model and then frozen, while only the dense parameters are updated.
Table 7. Online A/B test results based on a 30-day experimental window.
Metric GMV Orders Buyers CTR
Gains +7.03% +2.11% +1.86% +3.78%